OpenStudy (anonymous):

10.) system of equations r+s =-12 4r-6s=12

OpenStudy (anonymous):

Solve for r in the first equation r+s=-12 r=-12-s Substitute that expression into the next equation where r is 4(-12-s)-6s=12 now you have an equation with one variable, solve for s, then we can use that to solve for r.

OpenStudy (anonymous):

I got s=-6

OpenStudy (anonymous):

thanks?

OpenStudy (anonymous):

I got (-6,-6)

OpenStudy (anonymous):

is that right??

OpenStudy (anonymous):

looks right to me. =D

OpenStudy (anonymous):

thnx